# A Survey on 3D Gaussian Splatting

### v4

@@ 5.2 Dynamic Scene Reconstruction
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time [170, 171, 172, 173]. This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., virtual and augmented reality, 3D animation, and computer vision.
@@
To extend the concept of 3D GS to dynamic scenes, a straightforward approach is to incorporate the temporal dimension, allowing for the representation of scenes that change over time. 3D GS based methods [94, 93, 103, 174, 122, 175, 95, 118, 121, 120, 119, 176, 177, 123, 178] for dynamic scene reconstruction can generally be divided into two main categories. The first category utilizes additional structured information like spatial MLPs or grids (Sec. 4.6). For example, Yang et al. [94] first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS [125] involves the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. [118] introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. [93] designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes. While significant progress has been made in modeling dynamics and deformation at the Gaussian level, there is a compelling need to explore motion modeling at the object level, which might offer solutions to prevalent challenges like artifact reduction and the precise capture of fine-grained movements across long temporal sequences.
5.3 AI-Generated Content (AIGC)
AIGC refers to the digital content that is autonomously created or significantly altered by artificial intelligence systems, particularly in the fields of computer vision, natural language processing, and machine learning. AIGC is characterized by its ability to simulate, extend, or augment human-generated content, enabling applications that range from realistic image synthesis to dynamic narrative creation. The significance of AIGC lies in its transformative potential across various sectors, including entertainment, education, and technology development [179, 180, 181, 182]. It’s a pivotal element in the evolving landscape of digital content creation, offering scalable, customizable, and often more efficient alternatives to traditional methods.
This explicit nature of 3D GS facilitates real-time rendering capabilities and unprecedented levels of control and editability, making it highly relevant for and AIGC applications. The explicit scene representations and differentiable rendering algorithms of 3D GS align perfectly with the requirements of AIGC for generating high-fidelity, real-time, and editable content, crucial for applications in virtual reality, interactive media, and beyond. Recent works have effectively utilized 3D GS in conjunction with fields like generative models [183, 126, 127, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 128, 194, 195, 196, 129, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207], avatars [208, 209, 130, 210, 211, 212, 213, 214, 215, 131, 216, 132, 217, 218, 23, 219, 220, 221, 222, 223, 224, 225, 226, 227, 133, 228, 229], and scene editing [134, 135, 230, 91, 92, 90, 231, 121, 120, 119, 232, 233, 234, 235, 236, 237, 238, 136, 239, 102]. For instance, DreamGaussian [126] accelerated the generation of photorealistic 3D assets from single-view image through a three-step process: a diffusion based generative GS process, followed by an efficient algorithm for mesh extraction from 3D Gaussians based on local density querying, and finally a UV-space refinement stage to improve texture details. By combining 3D GS with a parametric morphable face model, GaussianAvatars [214] offered enhanced fidelity and flexibility in avatar animation, significantly improving upon current methods in novel-view rendering and expression reenactment. To improve the efficacy of text instruction based editing, Chen et al. [134] devised a semantic tracker to track the editing target during training while Fang et al. [135] proposed to extract the region of interest corresponding to the instruction. These advancements hold promise for numerous industrial applications, including digital assetization (with mesh extraction), the generation of long-form videos (e.g., Sora), etc.
5.4 Autonomous Driving
Autonomous driving aims to allow vehicles to navigate and operate without human intervention. These vehicles are equipped with a suite of sensors, including cameras, light detection and ranging (LiDAR), and radar, combined with advanced algorithms, machine learning models, and significant computational power [240, 241, 242, 243]. The central aim is to perceive the environment, make informed decisions, and execute maneuvers safely and efficiently [244, 245, 246, 247].
Autonomous vehicles need to perceive and interpret their surroundings to navigate safely. This involves reconstructing the driving scene in real-time, accurately identifying static and dynamic objects, and understanding their spatial relationships and movements [248, 249, 250]. In dynamic driving scenes, the environment is continuously changing due to moving objects like other vehicles, pedestrians, or animals [251]. Accurately reconstructing these scenes in real-time is crucial for safe navigation but is challenging due to the complexity and variability of the elements involved. In autonomous driving, 3D GS can be utilized to reconstruct a scene by blending data points (such as those obtained from sensors like LiDAR) into a cohesive and continuous representation. This is particularly useful for handling the varying densities of data points and ensuring a smooth and accurate reconstruction of both the static background and dynamic objects in a scene. To reconstruct complex 3D scenes from sparse sensor data, especially at high speeds and with moving objects, mainstream frameworks separated the urban/street scene into static and dynamic elements, where the dynamic elements are modeled using a composite dynamic Gaussian graph [137], point clouds combined with semantic logits [138], or physically constrained models [139]. By delving deeper into physics- and semantics-aware 3D GS (see “physics- and semantics-aware scene representation” in Sec. 7), one can expect that 3D GS could ideally serve as the cornerstone for environment perception in autonomous driving.
@@ 5.5 Endoscopic Scene Reconstruction
Surgical 3D reconstruction represents a fundamental task in robot-assisted minimally invasive surgery, aimed at enhancing intraoperative navigation, preoperative planning, and educational simulations through precise modeling of dynamic surgical scenes. Pioneering the integration of cutting-edge dynamic radiance fields into this domain, recent advancements have focused on surmounting the inherent challenges of single-viewpoint video reconstructions – such as occlusions by surgical instruments and sparse viewpoint diversity within the confined spaces of endoscopic exploration [252, 253, 254]. Despite the progress, the call for high fidelity in tissue deformability and topological variance remains, coupled with the pressing demand for faster rendering to bridge the utility in applications sensitive to latency [140, 141, 142]. This synthesis of immediacy and precision in reconstructing deformable tissues from endoscopic videos is essential in propelling robotic surgery towards reduced patient trauma and AR/VR applications, ultimately fostering a more intuitive surgical environment and nurturing the future of surgical automation and robotic proficiency.
Compared to typical dynamic scene reconstruction, endoscopic scene reconstruction faces unique challenges, e.g., sparse training data, due to constrained camera movement in confined space, unobserved regions due to tool occlusion, and pronounced non-rigid deformation of tissues. Existing approaches mainly used additional depth guidance to infer the geometry of tissues [140, 141, 142]. For instance, EndoGS [142] integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian [141] introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. [143] argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM [116] and Gaussian Pancake [144] devised SLAM systems for endoscopic scenes and showed significant speed advantages. Note that present efforts are mainly directed towards reconstructing from a single viewpoint, which continues to face obstacles in surgical applications.
@@ 5.6 Large-scale Scene Reconstruction
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy [146, 147, 148, 145] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. With the same spirit, Zhao et al. [152] proposed a distributed implementation of 3D GS training. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [44, 147]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [147], or designing explicit LoD structures like the Octree [148] and hierarchy [145]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [149, 150, 151]. One prominent challenge in large-scale scene reconstruction is limited capture, which could be alleviated by utilizing good priors (see “learning physical priors from large-scale data.” in Sec. 7).
6 Performance Comparison
TABLE I: Quantitative localization results6.1) on Replica [255], in terms of absolute trajectory error (ATE, cm). (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@ 6 Performance Comparison
HFSLAM [113] [IROS24] 0.19 0.34 0.16 0.21 0.26 0.23 0.21 0.38 0.25
TABLE II: Quantitative rendering results6.2) on Replica [255], in terms of PSNR, SSIM, and LPIPS. The numbers of FPS are taken from [106].

### v5

@@ 5.2 Dynamic Scene Reconstruction
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time [184, 185, 186, 187]. This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., VR/AR, 3D animation, and autonomous driving [188, 189, 190].
@@
The key to adapt 3D GS to dynamic scenes is the modeling of temporal dimension which allows for the representation of scenes that change over time. 3D GS based methods [94, 93, 106, 191, 129, 192, 95, 125, 128, 127, 126, 193, 194, 130, 195, 196, 197, 198, 199] for dynamic scene reconstruction can generally be divided into two main categories as discussed in Sec. 4.5 and Sec. 4.6. The first category utilizes additional fields like spatial MLPs or grids to model deformation (Sec. 4.6). For example, Yang et al. [94] first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS [132] devised the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. [125] introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. [93] designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
While 3D GS advances dynamic scene reconstruction by modeling per-Gaussian deformations, its reliance on fine-grained primitives limits scalability and robustness. Current methods struggle to balance computational efficiency and precision: small-scale reconstructions unify dynamic and static elements but become intractable in large environments, often requiring manual priors to segment regions —– a barrier in unstructured settings. Furthermore, the absence of object-level motion reasoning leads to artifacts and poor generalization over long sequences. Future work might prioritize object-centric frameworks that hierarchically group Gaussians into persistent entities, enabling efficient large-scale reconstruction through inherent motion disentanglement (dynamic vs. static).
5.3 Generation and Editing
Content generation and editing represent two fundamental and inherently interconnected capabilities in modern AI systems. While generation enables the synthesis of novel digital content from scratch or conditional inputs [200, 201, 202], editing provides the crucial ability to refine, adapt, and manipulate existing content with precise control [203]. Together, these capabilities revolutionize creative workflows by combining initial content creation with iterative refinement, enabling applications from professional content production to interactive consumer tools.
Recent advances in generation [204, 133, 134, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 135, 215, 216, 217, 136, 218, 219, 220, 221, 222, 223, 224, 225, 137, 226, 138, 227] have led to the emergence of three main approaches. Optimization based methods [204, 133, 134] distill diffusion priors (gradients) to guide 3D model updates with the score functions. While these methods demonstrate impressive fidelity, they face significant computational overhead due to the necessity of comparing multiple viewpoints during the optimization process. Reconstruction based methods [135, 227, 225] reframe the generation problem as a multi-view reconstruction task utilizing pre-trained multi-view diffusion models. Although this approach offers an intuitive and straightforward solution, it grapples with fundamental limitations in maintaining view consistency. The lack of strict geometric constraints across different viewpoints often results in inconsistent surface geometry and degraded texture quality, particularly in regions with complex visual features. Direct 3D generation methods train diffusion models on 3D representations [138, 220, 226]. While the learned 3D diffusion models facilitate multi-view consistency, the demanding computational costs impede the expansion of training scales necessary for improved generative diversity.
Current editing works [141, 142, 228, 91, 92, 90, 229, 128, 127, 126, 230, 231, 232, 233, 234, 235, 236, 143, 237, 140, 238, 239] fall into two primary classes. The first class leverages 2D image-editing models (e.g., diffusion-based editors) to iteratively refine 3D Gaussians. Early efforts [141, 142, 233] adopt optimization- or reconstruction-based strategies akin to methods in generation, but introduce task-specific control signals. However, naively applying 2D edits independently across views often introduces multi-view inconsistencies. Subsequent works [240, 140, 238, 239] mitigate this through iterative refinement or cross-view attention, albeit at increased computational costs for alignment. A notable challenge is unintended object deformations, attributed to the weak 3D geometric priors in 2D editing models and the difficulty of reconciling 2D edits with underlying 3D structures. The second class exploits the explicit nature of 3D GS to enable direct manipulation based on embedded properties such as semantics [232, 92, 91, 143] and key points [128]. However, this class remains underexplored due to essentail challenges: the lack of inherent ordering of Gaussians complicates the design of efficient indexing schemes, while editing attributes (e.g., texture and geometry) requires careful regularization and alignment to preserve plausibility.
5.4 Avatar
Avatars, the digital representations of users in virtual spaces, bridge physical and digital realms, enabling immersive interaction, identity expression, and remote collaboration. Spanning entertainment (gaming, virtual influencers), enterprise (AI agents, virtual meetings), healthcare, and education, they underpin metaverse economies. Advances in AR and VR amplify their role in redefining social, industrial, and creative landscapes.
3D GS has emerged as a powerful tool for human avatar reconstruction, primarily advancing along two directions: full-body modeling and head-centric modeling. For full-body avatars [241, 242, 144, 146, 243, 244, 245, 246, 145, 139, 247, 248, 249, 147, 250, 251, 252], the current methods typically anchor 3D Gaussians in a canonical space and deform them via parametric body models (e.g., SMPL) or cage-based rigging to model dynamic motions. These approaches adopt a hybrid deformation strategy: linear blend skinning handles rigid skeletal transformations such as joint rotations, while pose-conditioned deformation fields account for secondary non-rigid effects like muscle jiggles. For head avatars [149, 150, 253, 23, 254, 255, 148, 151, 256], the emphasis shifts to modeling intricate facial expressions, fine-grained geometry (e.g., wrinkles, hair [257]), and dynamic speech-driven animations. Techniques mainly combine parametric morphable face models (e.g., FLAME) with deformable 3D Gaussians, employing diffusion strategies and expression-aware deformation fields to disentangle rigid head poses from non-rigid facial movements. Both directions exploit the speed advantage and editability of 3D GS to enable efficient training, real-time rendering, and precise control over deformations, while addressing challenges in cross-frame correspondence, topology flexibility, and multi-view consistency.
Reconstruction in challenging scenes (e.g., occlusions, sparse single-view inputs, or loose clothing) and enhancing avatar interactivity represent critical challenges and opportunities. Parametric model-free methods, which bypass predefined priors by learning skinning weights directly from data, show promise for such scenarios. Complementary to this, generative models can mitigate ambiguities inherent in underconstrained settings. Further integrating physics-based constraints might bridge the gap between static reconstructions and responsive, lifelike interactions, unlocking applications in AR, embodied AI, etc.
@@ 5.5 Endoscopic Scene Reconstruction
Surgical 3D reconstruction represents a fundamental task in robot-assisted minimally invasive surgery, aimed at enhancing intraoperative navigation, preoperative planning, and educational simulations through precise modeling of dynamic surgical scenes. Pioneering the integration of dynamic radiance fields into this domain, recent advancements have focused on surmounting the inherent challenges of single-viewpoint video reconstructions such as occlusions by surgical instruments and sparse viewpoint diversity within the confined spaces of endoscopic exploration [258, 259, 260]. Despite the progress, the call for high fidelity in tissue deformability and topological variation remains, coupled with the pressing demand for faster rendering to bridge the utility in applications sensitive to latency [152, 153, 154]. This synthesis of immediacy and precision in reconstructing deformable tissues from endoscopic videos is essential in propelling robotic surgery towards reduced patient trauma and AR/VR applications, ultimately fostering a more intuitive surgical environment and nurturing the future of surgical automation and robotic proficiency.
Endoscopic scene reconstruction introduces distinct challenges compared to general dynamic scenes, including sparse training data from limited camera mobility in narrow cavities, frequent tool occlusions obscuring critical regions, and single-view geometry ambiguities. Existing approaches mainly used additional depth guidance to infer the geometry of tissues [152, 153, 154]. For instance, EndoGS [154] integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian [153] introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. [155] argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM [156] and Gaussian Pancake [157] devised SLAM systems for endoscopic scenes and showed significant speed advantages.
Advancing endoscopic 3D reconstruction requires targeted efforts in both data and dynamics modeling. Data limitations arise from single-viewpoint videos, which produce ill-posed reconstruction problems due to instrument occlusions and constrained camera mobility, leaving critical tissue regions unobserved. While depth estimators provide temporary workarounds, integrating multi-view camera systems addresses the root cause. In addition, existing datasets often feature truncated sequences (e.g., $4\!\sim\!8s$ in EndoNeRF [258]), which fail to capture prolonged tissue deformation dynamics or complex surgical workflows. Extending temporal coverage to include longer, clinically representative sequences would benefit downstream applications as aforementioned. Modeling limitations persist in current methods, which often represent tissue dynamics at the Gaussian level rather than object- or 3D region-level. This reduces their capacity to encode semantically meaningful anatomical interactions and deserves further explorations.
@@ 5.6 Large-scale Scene Reconstruction
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy [163, 164, 165, 162] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. With the same spirit, Zhao et al. [161] proposed a distributed implementation of 3D GS training. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [44, 164]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [164], or designing explicit LoD structures like the Octree [165] and hierarchy [162]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [158, 159, 160].
One prominent challenge in large-scale scene reconstruction lies in handling sparse or incomplete capture data, which can be mitigated through few-shot adaptation schemes (see Sec. 4.1) or generalizable priors (see “learning physical priors from large-scale data” in Sec. 7). Meanwhile, memory and computational bottlenecks can be addressed via distributed learning strategies [161], such as parameter partitioning across GPU clusters and parallel batched multi-view optimization.
TABLE I: Comparison of localization methods (§6.1) on Replica [261] (static scenes), in terms of absolute trajectory error (ATE, cm). (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@
5.7 Physics
The simulation of complex real-world dynamics, such as seed dispersal or fluid motion, is pivotal for applications spanning virtual reality, animation, and scientific modeling, where realism hinges on accurate physical behavior. Advances in video diffusion models have driven progress in 4D content generation, yet these methods might produce visually plausible results that violate fundamental physical laws. 3D GS emerges as a promising solution by embedding physical constraints and properties into scene representations, enabling both visually convincing and physically coherent simulations.
Existing methods differ in how they formulate and integrate physics-based priors into their frameworks. The most common approach is employing physics simulation engines (e.g., MLS-MPM [268]) to guide the dynamics generation. The material point method [268] and position based dynamics [269] — numerical methods used in computer graphics for simulating deformations in materials like fluids, granular media, and fracturing solids — have been extensively explored by the community through various customizations [166, 167, 168, 143, 169, 170, 21, 171]. Analytical material models, such as mass-spring systems, have also demonstrated success in approximating deformations by explicitly encoding material properties into 3D Gaussians [172]. Across these methods, 3D Gaussians are treated as discrete particles (with one exception [173] using a continuous representation) and serve as computational units within the chosen simulator. Unknown material properties or physical parameters are typically learned through video-based supervision from conditional generative models.
Despite advancements in physics based 3D GS frameworks, critical limitations persist. Current systems struggle to unify diverse physical behaviors (e.g., rigid, elastic, or soft-body dynamics) into cohesive simulations, handle complex multi-object interactions without manual intervention, and model scene-level interactions such as environmental feedback and dynamic lighting changes. Integrating adaptive physics engines capable of multi-object and multi-material interactions, developing new simulation architectures that are compatible with priors learned from large-scale data, and expanding datasets to encompass diverse materials and dynamic scenarios are equally vital.
6 Performance Comparison
TABLE II: Collection of representative datasets for 3D GS. Here PC represents point clouds.
       Name Type # Sample Task
Tanks&Temples [270] [TOG17] RGB 14 Novel View Synthesis
RealEstate10K [271] [TOG18] RGB 1,000
DeepBlending [272] [TOG18] RGB 19
LLFF [273] [TOG19] RGB 8
NeRF [12] [ECCV20] RGB 8
ACID [274] [ICCV21] RGB 700+
Mip-NeRF 360 [8] [CVPR22] RGB 9
TUM RGB-D [275] [IROS12] RGB-D 39 Robotics
KITTI [276] [CVPR12] RGB-D&PC 11
ScanNet [277] [CVPR17] RGB-D 1,513
Replica [261] [arXiv19] RGB-D 18
Waymo [278] [CVPR20] RGB-D&PC 1,150
nuScenes [279] [CVPR20] RGB-D&PC 1,000
RLBench [280] [RA-L20] RGB 100
Robomimic [281] [CoRL22] RGB 800
D-NeRF [184] [CVPR21] RGB 8 Dynamic Scene Reconstruction
HyperNeRF [185] [TOG21] RGB 6
NeRF-DS [282] [CVPR23] RGB 8
CoNeRF [283] [CVPR22] RGB 7 Generation and Editing
SPIn-NeRF [284] [CVPR23] RGB 10
Tensor4D [285] [CVPR23] RGB 4
OmniObject3D [286] [CVPR23] 3D Object 6,000
Objaverse [287] [CVPR23] 3D Object 800K+
People-Snapshot [288] [CVPR18] RGB 24 Avatar
VOCASET [289] [CVPR19] RGB 12
THUman [290] [ICCV19] RGB 200
THUman2.0 [291] [CVPR21] RGB-D 500
ZJU-Mocap [292] [CVPR21] RGB 9
H3DS [293] [ICCV21] RGB 23
THUman3.0 [294] [TPAMI22] 3D Scan 20
SCARED [295] [MICCAI19] RGB-D 9 Medical
EndoNeRF [258] [MICCAI22] RGB 2
X3D [296] [CVPR24] X-ray 15
CityNeRF [297] [ECCV22] RGB 12 Large-scale Reconstrction
Waymo Block-NeRF [298] [CVPR22] RGB&PC 1
UrbanBIS [299] [SIGGR23] RGB&PC 6
GauU-Scene [160] [arXiv24] RGB&PC 1
TABLE III: Comparison of mapping methods (§6.2) on Replica [261] (static scenes), in terms of PSNR, SSIM, and LPIPS. The results for FPS are taken from [113] using one 4090 GPU.